OpenStudy (anonymous):

Let the function f be defined by y = f(x), where x and f(x) are real numbers. f(x)=8 over x-4 f(2) f(-3) f(k) f(k^2-1)=

OpenStudy (amistre64):

its really just a matter of plugging in those f numbers in place of the x

OpenStudy (amistre64):

\[f(2)=\frac{8}{2-4}\] \[f(-3)=\frac{8}{-3-4}\] \[f(k)=\frac{8}{k-4}\] \[f(k^2 -1)=\frac{8}{k^2 -1-4}\]

OpenStudy (anonymous):

sweet.. that is what i came up with on this one.. and i beleive if i simplify the f(2)= 8 over2-4 it comes out to -4... is that correct?

OpenStudy (amistre64):

yes, thats right